# MeanCache: User-Centric Semantic Cache for Large Language Model Based Web Services

### v3

@@ Abstract.
Large Language Models (LLMs) like ChatGPT and Llama have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters, where inference demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries, which constitute about 31% of the total queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries nor do they operate on contextual queries, leading to unacceptable false hit-and-miss rates.
@@
This paper introduces MeanCache, a user-centric semantic cache for LLM-based services that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model without violating user privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency and costs and enhances model performance, resulting in lower false hit rates. MeanCache also encodes context chains for every cached query, offering a simple yet highly effective mechanism to discern contextual query responses from standalone. Our experiments benchmarked against the state-of-the-art caching method, reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions while performing even better on contextual queries. It also reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%.
Federated Learning, Caching, Large Language Models, Cache, Hit, Miss, Rate, ChatGPT, Llama, Semantic, Search, Large Language Models, LLMs, GPT, Machine Learning, Privacy, Web, Embeddings, Cosine Similarity, Client, Prompt, Query,
1. Introduction
@@
Large Language Models (LLMs) like ChatGPT (cha, [n. d.]), Google Bard (bar, [n. d.]), Claude (cla, [n. d.]), and Llama 2 (Touvron et al., 2023) have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in applications ranging from search engines to conversational agents. LLMs are increasingly integrated into platforms like the Perplexity AI search engine, Rabbit OS (rab, [n. d.]), Arc browser (arc, [n. d.]), and Bing Chat (bin, [n. d.]). For instance, in 2023, 1.9 billion queries were sent to Bing chat, reflecting the massive use of such systems (bin, [n. d.]).
@@
Motivation. Generating responses to user queries with LLMs, such as GPT-3, requires substantial computations and poses environmental challenges (Frantar et al., 2023; Wang et al., 2023; Tseng et al., 2024; Sachdeva et al., 2024; Ma et al., 2024; Strubell et al., 2019). For example, GPT-3’s 175B parameters in float16 format consume 326 GB memory, exceeding single GPU capacities and necessitating multi-GPU deployments (Frantar et al., 2023). These requirements lead to high operational costs.
Consequently, LLM-based services charge users and limit query rates (per, [n. d.]; ope, [n. d.]). Prior studies report that users frequently submit similar queries to web services (Lempel and Moran, 2003; Xie and O’Hallaron, 2002)—approximately 33% of search engine queries being resubmitted (Markatos, 2001). We investigate if a similar pattern exists with queries to LLMs. Our analysis of a large corpus of real-world user conversations with ChatGPT reveals that approximately 31% of such queries have a high similarity with a previously submitted query, as shown in Section 2. Thus, there is a potential to reduce the query inference cost of LLMs by up to one-third.
@@
Caching serves as an effective technique in traditional web services to address duplicate search queries, avoiding redundant computations, significantly improving response time, reducing the load on query processors, and enhancing network bandwidth utilization (Markatos, 2001; Saraiva et al., 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003; Brin and Page, 1998; Podlipnig and Böszörmenyi, 2003; Baeza-Yates et al., 2007). If applicable to LLMs-based web services, such caching can substantially impact billions of floating point operations, thereby decreasing operational costs and environmental impacts.
Figure 1. Analysis of real-world ChatGPT conversations reveals that, on average, 31% of queries are similar to previously submitted queries.
@@
Problem. Existing caching techniques  (Saraiva et al., 2001; Markatos, 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003) use keyword matching, which often struggles to capture the semantic similarity among similar queries to LLM-based web services, resulting in a significantly low hit rate. For instance, existing caches do not detect the semantic similarity between ”How can I increase the battery life of my smartphone?” and ”Tips for extending the duration of my phone’s power source.”, leading to a cache miss. Recently, Zhu et al. (2023) and GPTCache(Bang, 2023) present server-side semantic caching for the LLMs-based services to address the limitations of keyword-matching caching techniques. If a new query is semantically similar to any query in the cache, the server returns the response from the cache. Otherwise, a model multiplexer selects the most suitable LLM for the query to generate the response.
@@
Existing semantic caches have several limitations. First, they demand a significantly large central cache to store the queries and responses of all users, which is unscalable and violates users’ privacy. Second, they incur the network cost of sending a user query to the server even if there is a cache hit. An end user will still be charged for the query even if the query is served from the server cache. Third, they use a single server-side embedding model to find the semantic similarity among queries, which does not generalize to each user’s querying patterns. For instance, Google Keyboard (gbo, [n. d.]) adapts to each user’s unique writing style and embeds such personalized behaviors to enhance the accuracy of the next word prediction model. Fourth, they employ Llama-2 to enhance the accuracy of semantic matching (Bang, 2023); however, in practice, such models perform billions of operations to generate embeddings, offsetting the benefits of the cache. Lastly, they are only effective on standalone LLM queries, resulting in unbearably high false hit rates for contextually different but semantically similar queries.
@@
To address privacy concerns associated with central server-side caching, MeanCache introduces a user-side cache design ensuring that the user’s queries and responses are never stored outside of the user’s device. While a personalized per-user cache may not offer benefits for similar queries across multiple users, our investigation with real-world conversations with ChatGPT reveals that approximately 31% of the queries are still similar to at least one of the previously submitted queries by the same user. Therefore, having a user-centric cache offers the same, if not more, opportunities to eliminate expensive LLM queries.
@@
To find a semantic match between a new query and cached queries, MeanCache uses smaller embedding models such as MPNet (Song et al., 2020) to generate embeddings for semantic matching locally. Previous work has shown that a smaller model can achieve performance comparable to larger models on custom tasks (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). Prior work (Bang, 2023) proposes using large language models, such as Llama 2, to generate embeddings. However, these models require billions of operations, making them unsuitable for embeddings generation.
@@
Due to different contexts around queries, LLM may return different responses for semantically similar queries. For such queries, MeanCache also records the contextual chain, parent queries already in the cache, for a given query. To find a response for a contextual query, MeanCache verifies the context of a contextual query by matching a given query’s context with the cached query’s context chain to accurately retrieve responses for previously seen contextually different but semantically similar queries from the cache.
@@
Each user may not have sufficient queries to customize an embedding model that can help find a semantic match between new queries and cached queries. To address this, MeanCache utilizes federated learning, which exploits data silos on user devices for private training for collaborative learning, thereby personalizing an embedding model for each user. This privacy-preserving training not only customizes the embedding model to the user’s querying patterns but also enhances the performance (i.e., accuracy) of semantic caching without compromising user privacy (i.e., without storing user data on the web server).
@@
The runtime performance of MeanCache is primarily influenced by the time taken to match a new query embedding vector with existing ones in the cache to find a semantically similar query. The search time is directly proportional to the dimensions of the embedding vector. To optimize runtime performance, MeanCache compresses the embedding vector by leveraging principal component analysis (PCA) (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), effectively reducing the size of the embedding vector (i.e., projecting it to lower dimensional space). MeanCache also offers an adaptive cosine similarity threshold, which is also collaboratively computed using FL, to improve accuracy in finding semantic matches between queries.
Study of Real-World Queries. To measure the prevalence of similar queries in real-world conversation with LLM-based web service, We recruited 20 participants who regularly use ChatGPT and conducted a thorough analysis of their queries in a privacy-preserving manner. Across 27K queries, we find that 31% of a user’s queries are similar to at least one previous query by the same user, which could have been retrieved from a cache.
@@
Evaluations. We compare MeanCache with GPTCache (Bang, 2023), a widely-used open-source semantic cache for LLM-based web services. GPTCache (Bang, 2023) is the most closely related work to MeanCache and has received over 6,000 stars on GitHub (zil, [n. d.]). We benchmark MeanCache’s performance against GPTCache on the GPTCache’s dataset (dat, [n. d.]) to demonstrate its effectiveness and highlight the improvements MeanCache offers over existing solutions.
@@
MeanCache surpasses GPTCache (Bang, 2023) by achieving a 17% higher F-score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s performance on contextual queries is even more impressive when compared to GPTCache (baseline). For contextual queries, MeanCache achieves a 25% higher F-score and accuracy, and a 32% higher precision over the baseline. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching while still outperforming the state-of-the-art GPTCache.
@@
Artifact Availability: MeanCache is implemented in the Flower FL framework (Beutel et al., 2020). The complete source code and contextual queries dataset will be available upon acceptance.
2. Similar User Queries in Practice
Prior study (Markatos, 2001) measured the prevalence of similar queries across the entire user population. To understand the extent to which a single user submits similar queries, we conducted the first empirical study to measure how often a real-world ChatGPT user submits similar queries. Given the private nature of conversations with the ChatGPT service, study participants are naturally reluctant to share their query corpus. To address this problem, we develop a script that runs locally on each participant’s computer and identifies semantically similar queries, and calculates the percentage of similar queries on the participant’s query corpus. We then ask the participants to review the results and share, if comfortable, only three metrics (of similar queries, total queries, and days since the first query) anonymously with the authors.
We recruited 20 participants, including male and female participants, university professors, and graduate students who regularly use ChatGPT. Before measuring the query statistics, our analysis script addresses false positives by excluding large queries with lengths beyond 256 tokens. We examined over 27K queries across 20 participants who have been using ChatGPT for an average of 332 days.
Figure 1 presents the results of this study. We find that, on average, 31% of each participant’s queries are highly similar to at least one of the previously submitted queries by the same participant, which necessitates a user-centric cache for LLM-based web services.
3. Background
3.1. Federated Learning
@@
Federated Learning is a distributed, privacy-preserving training of a machine learning (ML) model. An FL round starts when a central server shares a global model with all the participating clients. Each client trains the model on its local data and sends the updated model to the central server. The central server aggregates the received models’ weights to form a new global model to be used in the next round. Different aggregating algorithms have been proposed, such as FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), and FedMA (Wang et al., 2020b). The FedAvg algorithm is the most popular one, and it aggregates the models with the following equation:
(1) $W_{global}^{t+1}=\sum_{k=1}^{K}\frac{n_{k}}{n}w_{k,t}$
@@
$W_{global}^{t+1}$represents the new global model, $w_{k,t}$ denotes the model of the $k^{th}$ client at the $t^{th}$ round, $n_{k}$ are the number of samples available at $k^{th}$ client and $n$ is the total number of samples of clients participating in the given round. This procedure is reiterated for multiple rounds until the model reaches convergence.
3.2. Transformer Architecture
The transformer architecture is a deep learning model based on the attention mechanism, which allows the model to focus on the important parts of the input (Vaswani et al., 2017). The transformer architecture is primarily used for sequence-to-sequence tasks in NLP, such as machine translation, text summarization, and question-answering.
The attention mechanism calculates the importance of each input element by considering its relationship with other elements in the sequence. This enables the model to capture long-range dependencies and improve performance on tasks that require an understanding of context. The feed-forward layer helps the model capture complex patterns and relationships in the data. The transformer architecture also incorporates positional encoding, which provides information about the position of each element in the input sequence. This helps the model understand the order of the elements, which is crucial for tasks like machine translation.
3.2.1. Embeddings and Cosine Similarity
Transformer model can encode the given text into embeddings. Embeddings represent words or phrases as dense vectors in a high-dimensional space and capture the semantic meaning of words or phrases, and therefore, words or sentences with similar meanings will have similar vector representations (Reimers and Gurevych, 2019; Karpukhin et al., 2020; Ni et al., 2022). For example, the words ”king” and ”queen” are semantically similar because they both refer to royalty. Therefore, their embeddings will be close to each other in the vector space. These embeddings encode the contextual information and semantic meaning of the sentences. Cosine similarity is a measure of similarity between two vectors. In the context of embeddings, cosine similarity is often used to quantify the similarity between sentence embeddings. It is calculated by taking the cosine of the angle between the two vectors. The cosine similarity ranges from -1 to 1, where -1 indicates completely opposite vectors, 0 indicates orthogonal (perpendicular) vectors, and 1 indicates identical vectors. Cosine similarity ($\theta$) between embeddings E1 and E2 is measured as:
(2) $\theta=\frac{{\textbf{E1}}\cdot{\textbf{E2}}}{\|{\textbf{E1}}\|\|{\textbf{E2}}\|}$
Here, E1 and E2 represent the two word embeddings, ”dot” represents the dot product of the vectors, and ”——E1——” and ”——E2——” represent the Euclidean norms (magnitudes) of the vectors.
Figure 2. MeanCache’s Workflow
4. MeanCache’s Design
@@
MeanCache is a user-centric semantic caching system optimized for user-side operation. Figure 2 illustrates the workflow of MeanCache for similar queries. Algorithm 1 further explains MeanCache’s querying and population process programmatically. When a user submits a query to an LLM web service with MeanCache enabled, MeanCache computes the query embeddings (Line  3). These embeddings are then matched with the embedding of the cached queries using cosine similarity (Line  4). For every similar query found within the cache, MeanCache analyzes the context chain for every query (Line  7) and matches its embedding with the conversational history (Line  9) of the submitted query. If MeanCache finds a similar query with a similar context chain, the response is retrieved from the local cache and returned to the user (Line  13). Otherwise, MeanCache forwards the query to the LLM web service to obtain the response. The query and its response and embeddings are then stored in the cache (Line  18).
@@
The key contribution of MeanCache comes from harnessing the collective intelligence of multiple users to training a semantic similarity model and its user-centric design that addresses both privacy and scalability issues. To achieve these, MeanCache takes the following design decisions. It employs a small embedding model with lower computational overhead than large Language Learning Model (LLM) based embedding models, such as Llama 2. It uses federated learning (FL) for collaborative training to fine-tune the embedding model without ever storing user data on a central server. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. To handle contextual queries, MeanCache includes contextual chain information in its cache against every query to identify if the cached response for a query is only applicable under a specific context. This design is capable of handling long contextual chains. To reduce storage and memory overhead and expedite the search time for finding similar queries in the cache, MeanCache compresses the embeddings using Principal Component Analysis (PCA). MeanCache leverages user feedback (i.e., forcing LLM request instead of the cache query) to adapt the threshold for cosine similarity. It then shares the updated threshold via FL to determine the optimal threshold across multiple users. It does this by varying the threshold and selecting the one that optimizes the F-score of the cache on clients.
Algorithm 1 MeanCache Workflow for LLM Queries
1:Input: User query $Q$, User Query Context $C_{q}$
2:Output: Response $R$
3:Compute query embedding $E_{Q}$
4:$similar\_queries\leftarrow$ FindSimilarQueriesinCache($E_{Q}$)
@@
5:$context\_match\leftarrow$ False
6:for each query $C_{i}$ in $similar\_queries$ do
7:    $ctx\leftarrow$ the entire context chain of $C_{i}$
8:    $E_{ctx}\leftarrow$ Retrieve embedding of $ctx$
9:    $E_{cq}\leftarrow$ Retrieve embedding of $C_{q}$
10:    Compute cosine similarity $s_{i}$ between $E_{cq}$ and $E_{ctx}$
11:    if $s_{i}\geq$ context similarity threshold then
12:        $context\_match\leftarrow$ True
13:        Retrieve response $R$ from cache for $C_{i}$
14:        return $R$
15:    end if
16:end for
17:if not $context\_match$ then
18:    $R\leftarrow$ LLMResponseAndEnrollInCache($Q$, $E_{Q}$ , $C_{q}$)
19:end if
20:return $R$
Figure 3. Privacy-Preserving Embedding Model FL Training in MeanCache.
4.1. FL Based Embedding Model Training

### v4

# MeanCache: User-Centric Semantic Caching for LLM Web Services

@@ Abstract
Large Language Models (LLMs) like ChatGPT and Llama have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters, where inference demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries nor do they operate effectively on contextual queries, leading to unacceptable false hit-and-miss rates.
@@
This paper introduces MeanCache, a user-centric semantic cache for LLM-based services that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model without violating user privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency, costs, and enhances model performance, resulting in lower false-hit rates. MeanCache also encodes context chains for every cached query, offering a simple yet highly effective mechanism to discern contextual query responses from standalone queries. Our experiments benchmarked against the state-of-the-art caching method reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions while performing even better on contextual queries. It also reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%.
Index Terms:
Large Language Models, Semantic Cache, Embedding, Contextual Queries, Cache, Privacy-Preserving AI
I Introduction
@@
Large Language Models (LLMs) like ChatGPT [1], Google Bard [2], Claude [3], and Llama [4] have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in applications ranging from search engines to conversational agents. LLMs are increasingly integrated into platforms like the Perplexity AI search engine, Rabbit OS [5], and Arc browser [6].
@@
Motivation. Generating responses to user queries with LLMs, such as GPT-3, requires substantial computations and poses environmental challenges [7, 8, 9, 10]. For example, GPT-3’s 175B parameters in float16 format consume 326 GB memory, exceeding single GPU capacities and necessitating multi-GPU deployments [7]. These requirements lead to high operational costs. Consequently, LLM-based services charge users and limit query rates [11, 12]. Prior studies have observed that users frequently submit similar queries to web services [13, 14, 15] (approximately 33% of search engine queries being resubmitted [15]), suggesting opportunities for optimization by avoiding redundant computations.
@@
Caching serves as an effective technique in traditional web services to address duplicate search queries, avoiding redundant computations, significantly improving response time, reducing the load on query processors, and enhancing network bandwidth utilization [15, 13, 16, 17]. If applicable to LLMs-based web services, such caching can substantially impact billions of floating point operations, thereby decreasing operational costs and environmental impact.
@@
Problem. Existing caching techniques  [18, 15, 14, 13] use keyword matching, which often struggles to capture the semantic similarity among similar queries to LLM-based web services, resulting in a significantly low hit rate. For instance, existing caches do not detect the semantic similarity between “How can I increase the battery life of my smartphone?” and “Tips for extending the duration of my phone’s power source”, leading to a cache miss. Recently, Zhu et al. [19] and GPTCache [20] present server-side semantic caching for the LLMs-based services to address the limitations of keyword-matching caching techniques. If a new query is semantically similar to any query in a cache, the server returns the response from the cache. Otherwise, a model multiplexer selects the most suitable LLM for the query to generate the response.
@@
Existing semantic caches have several limitations. First, they demand a significantly large central cache to store the queries and responses of all users, which is unscalable and violates users’ privacy. Second, they incur the network cost of sending a user query to the server even if there is a cache hit. An end user will still be charged for the query even if the query is served from the server cache. Third, they use a single server-side embedding model to find the semantic similarity among queries, which does not generalize to each user’s querying patterns. For instance, Google Keyboard [21] adapts to each user’s unique writing style and embeds such personalized behaviors to enhance the accuracy of the next word prediction model. Fourth, they employ Llama-2 to enhance the accuracy of semantic matching [20]; however, in practice, such models perform billions of operations to generate embeddings, offsetting the benefits of the cache. Lastly, they are only effective on standalone queries, resulting in unbearably high false hit rates for contextually different but semantically similar queries.
@@
To address privacy concerns associated with central server-side caching, MeanCache introduces a user-side cache design ensuring that the user’s queries and responses are never stored outside of the user’s device. To find a semantic match between a new query and cached queries, MeanCache uses smaller embedding models such as MPNet [22] to generate embeddings for semantic matching locally. Previous work has shown that a smaller model can achieve performance comparable to larger models on custom tasks [23, 24, 25].
@@
Due to different contexts around queries, LLM may return different responses for semantically similar queries. For such queries, MeanCache also records the contextual chain, parent queries already in the cache, for a given query. To find a response for a contextual query, MeanCache verifies the context of a contextual query by matching a given query’s context with the cached query’s context chain to accurately retrieve responses to contextual queries.
@@
Each user may not have sufficient queries to customize an embedding model that can help find a semantic match between new queries and cached queries. To address this, MeanCache utilizes federated learning (FL), which exploits data silos on user devices for private training for collaborative learning [26, 27, 28, 29], thereby personalizing an embedding model for each user. This privacy-preserving training not only customizes the embedding model to the user’s querying patterns but also enhances the performance (i.e., accuracy) of semantic caching without compromising user privacy (i.e., without storing user data on the web server).
@@
The runtime performance of MeanCache is primarily influenced by the time taken to match a new query embedding vector with existing ones in the cache to find a semantically similar query. The search time is directly proportional to the dimensions of the embedding vector. To optimize runtime performance, MeanCache compresses the embedding vector by leveraging principal component analysis (PCA) [30, 31, 32], effectively reducing the size of the embedding vector (i.e., projecting it to lower dimensional space). MeanCache also offers an adaptive cosine similarity threshold, which is also collaboratively computed using FL, to improve accuracy in finding semantic matches between queries.
@@
Evaluations. We compare MeanCache with GPTCache [20]. GPTCache is closely related to MeanCache and has received over 6,000 stars on GitHub [33]. We benchmark MeanCache’s performance against GPTCache on the GPTCache’s dataset [34] to demonstrate its effectiveness and highlight the improvements MeanCache offers over existing solutions.
@@
MeanCache surpasses GPTCache [20] by achieving a 17% higher F-score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s performance on contextual queries is even more impressive when compared to GPTCache (baseline). For contextual queries, MeanCache achieves a 25% higher F-score and accuracy, and a 32% higher precision over the baseline. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching while still outperforming the state-of-the-art GPTCache.
@@
Artifact Availability: MeanCache is implemented in the Flower FL framework [35]. The complete source code and contextual queries dataset will be available at https://github.com/SEED-VT/MeanCache.
II Background
@@
Federated Learning (FL). FL is a distributed, privacy-preserving ML model training approach [36, 27, 37, 38]. In each FL round, a central server distributes a global model to participating clients, who train it on their local data and return the updated models. The server then aggregates these models to create a new global model for the next round. While several aggregation algorithms exist (FedAvg [28], FedProx [39], and FedMA [40]), FedAvg is most common, using the equation:
$W_{global}^{t+1}=\sum_{k=1}^{K}\frac{n_{k}}{n}w_{k,t}$ (1)
@@
where $W_{global}^{t+1}$ represents the new global model, $w_{k,t}$ is the model of the $k^{th}$ client at round $t$, $n_{k}$ is the sample count at the $k^{th}$ client, and $n$ is the total sample count across participating clients. This process continues for multiple rounds until convergence.
Transformer and Embeddings. The transformer architecture is based on the attention mechanism [41]. Transformers are extensively utilized in natural language processing (NLP) tasks such as machine translation, text summarization, and question-answering. The attention mechanism allows transformers to effectively capture long-range dependencies within sequences, while positional encoding explicitly incorporates information about the order of tokens. Transformers convert text into embeddings, representing words or sentences as dense vectors in high-dimensional space. These embeddings capture semantic meanings, ensuring that semantically similar words or sentences have similar vector representations [42, 43, 44]. Cosine similarity is a widely used metric for measuring the similarity between embeddings, with values ranging from -1 (completely opposite) to 1 (identical). For instance, embeddings for words such as cat and dog typically exhibit higher cosine similarity compared to cat and car, as the former pair is more closely related semantically. For embedding vectors $\mathbf{E}_{1}$ and $\mathbf{E}_{2}$, cosine similarity is calculated as:
[…]
where $\mathbf{E}_{1}\cdot\mathbf{E}_{2}$ represents the dot product and $||\mathbf{E}_{1}||$ and $||\mathbf{E}_{2}||$ denote the vector magnitudes.
Contextual Queries. Interaction between end-users and LLM services typically involves two primary types of queries: standalone queries (e.g., Q1 Draw a line in Python?) and follow-up, or contextual queries (e.g., Q2 Change the color to red). Standalone queries can be resolved independently without any additional information, whereas contextual queries require prior context to provide accurate and relevant responses.
Consider a scenario in which both queries Q1 and Q2, along with their responses, have been cached. If a user later submits a new standalone query, Q3 Draw a circle?, followed by a contextual query, Q4 Change the color to red, Q4 might appear semantically similar to the cached query Q2. However, despite their similarity, Q4 references a different context (Q3 rather than Q1), thus requiring a distinct interpretation and response. Consequently, a semantic cache without effective context detection may yield incorrect cache hits.
Figure 1: MeanCache’s Workflow.
III MeanCache’s Design
@@
MeanCache is a user-centric semantic cache optimized for user-side operation. Figure 1 illustrates the workflow of MeanCache for similar queries. Algorithm 1 further explains MeanCache’s querying and population process programmatically. When a user submits a query to an LLM web service with MeanCache enabled, MeanCache computes the query embeddings (Line 1). These embeddings are then matched with the embedding of the cached queries using cosine similarity (Line 1). For every similar query found within the cache, MeanCache analyzes the context chain for every query and matches it with the conversational history of the submitted query (Line 1). If MeanCache finds a similar query with a similar context chain, the response is retrieved from the local cache and returned to the user (Line 1). Otherwise, MeanCache forwards the query to the LLM service to obtain the response. The query, its response, and embeddings are then stored in the cache (Line 1).
@@
MeanCache harnesses the collective intelligence of multiple users to train a semantic similarity model, and its user-centric design addresses privacy and scalability issues. To achieve these, MeanCache takes the following design decisions. It employs a small embedding model with lower computational overhead than LLM based embedding models. It uses FL for collaborative training to fine-tune the embedding model without ever storing user data on a central server. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. To handle contextual queries, MeanCache includes contextual chain information in its cache against every query to identify if the cached response for a query is only applicable under a specific context. This design is capable of handling contextual chains. To reduce storage and memory overhead and expedite search time for finding similar queries in the cache, MeanCache compresses the embeddings using PCA.
1
Input: User query $Q$, User Query Context $C_{q}$ (e.g., $null$ if no parent)
Output: Response $R$
2
$E_{Q}\leftarrow\text{encode}(Q)$ // compute the embedding of the query
3
$similar\_queries\leftarrow\text{FindSimilarQueriesinCache}(E_{Q})$ // retrieve top-$k$ similar cached queries
4
@@
$context\_match\leftarrow\textbf{False}$ // flag to indicate if suitable context is found
5
6foreach context $C_{i}\in similar\_queries$ do
7
8      if $C_{i}$ matches with $C_{q}$ then
9             $context\_match\leftarrow\textbf{True}$
10             Retrieve response $R$ from cache for $C_{i}$
11
12
13if not $context\_match$ then
       $R\leftarrow\text{LLMResponseAndEnrollInCache}(Q,E_{Q},C_{q})$ // generate response and cache it
14
15
16return $R$
Algorithm 1 MeanCache
III-A FL Based Embedding Model Training